if demand is 75 lakh setup and holding cost is 29 and 35 number of piece per order is 2.3 lakh printing is done 300 days and delivery within 2 days what is eoq?


1
Expert's answer
2021-07-09T12:24:01-0400

Answer:


Given that

Annual demand = 75 lakhs per year

Setup cost = $29 per unit

Holding cost = $35 per unit

Number of pieces in one order = 2.3 lakhs

Working days = 300 days

Lead time = 2 days

The formula to calculate EOQ is


EOD= "\sqrt{2SD \over H}"


Here, S = order cost

D = annual demand

H = holding cost

Substituting the values,


EOD= "\sqrt{2 \times 29 \times7500000 \over 35}"


EOD= "\sqrt{12428571.42}"


EOD= 3525.41